### v2

# Topological Center of the Double Dual of the Orlicz Figà-Talamanca Herz Algebra

Abstract.
Let a locally compact group and be a complementary pair of Young functions. Let be the Orlicz analogue of the classical Figà-Talamanca Herz algebra In this article, we establish a necessary and sufficient condition for the equality to hold, where denotes the topological center of the double dual of when equipped with the first Arens product. Furthermore, we prove several results concerning the semi-simplicity of the Banach algebras and
Key words and phrases:
Amenability; Arens product; Banach algebras; Figà-Talamanca Herz algebras; Orlicz spaces; Topological center.Corresponding author.
Email address: arvishdabra3@gmail.com (A. Dabra).
2020 Mathematics Subject Classification:
Primary 43A15, 46J10, 46E30; Secondary 43A07, 43A251. Introduction
Let be a Banach algebra and let denote its double dual. In 1951, Arens [4] introduced the first and second Arens products on proving that is a Banach algebra when equipped with either of these products. In this article, we refer to the first Arens product as the Arens product and denote it by From the definition, it follows that the map is continuous on for any In contrast, the map is not necessarily continuous. This asymmetry motivates the definition of the topological center of given by
It is straightforward to verify that
Let be the norm closure of the linear span of in The dual space also becomes a Banach algebra when equipped with the multiplication induced by the Arens product on Over the past three to four decades, the study of the topological center and related aspects of the algebras and has been the subject of considerable research interest. In 1987, Isik et al. [30] addressed the topological center problems for the group algebra and established that the topological center of the algebra coincides with for any compact group
Earlier, in 1964, Eymard [16] introduced and studied the Fourier algebra and its dual for any locally compact group If the underlying group is abelian with as its dual group, then the Fourier algebra is isometrically isomorphic to Later, in 1974, Granirer [24] defined the subspace of uniformly continuous functionals on denoted by When is abelian, this space coincides with the space of bounded uniformly continuous valued functions on the dual group
Cecchini and Zappa [6] as well as Lau and Losert [36, 37], investigated the topological center problems for the Banach algebra and the corresponding space Their results show that, for a large class of locally compact groups including the abelian groups, discrete amenable groups, the motion group, the affine group and the Heisenberg group, the topological center of the algebra is and the topological center of is the Fourier-Stieltjes algebra In 1996, Lau and Ülger [39] extended this line of research further analyzing the topological center problems for both the group algebra and the Fourier algebra yielding significant results. In 2016, Losert [40] proved that if contains a free group then the topological center of is strictly larger than
For and any locally compact group Figà-Talamanca [17] and Herz [28] introduced the -analogue of the Fourier algebra known as the Figà-Talamanca Herz algebra, denoted by The algebra of pseudomeasures, denoted is defined as the closure of in where is the representation of on given by left convolution. It is proved that the dual space of is isometrically isomorphic to where and are dual Hölder exponents. In the special case the algebra and its dual coincide with the Fourier algebra and the group von Neumann algebra respectively.
The algebra has received significant attention in recent decades from numerous researchers, including Daws [11], Derighetti [13, 14], Forrest [18, 19], Granirer [24, 25, 26, 27], Lau and Ülger [38] and Miao [41, 42, 43, 44], among others. The corresponding algebra of pseudomeasures has also been the subject of extensive study, notably in the work of Daws and Spronk [12] as well as Gardella and Thiel [20, 21, 22, 23]. In 2009, Miao [44] investigated the topological center problems for the algebra and the corresponding space establishing a necessary and sufficient condition for the equality to hold.
It is well known that the Orlicz spaces are an important generalization of the classical Lebesgue spaces and have attracted considerable attention. Recently, Lal and Kumar [31] and independently Aghababa and Akbarbaglu [1] introduced and studied the Orlicz -analogue of the Figà-Talamanca Herz algebra by replacing the spaces with spaces, where is a Young function satisfying the condition. The resulting algebra, denoted is known as the Orlicz Figà-Talamanca Herz algebra. If is a complementary pair of Young functions, then the dual space of is isometrically isomorphic to the space of pseudomeasures, denoted by [31, Theorem 3.5]. Both the algebra and its dual have received a great deal of interest in recent years; see [1, 2, 3, 7, 8, 9, 31, 32, 33, 34].
In this article, we investigate the topological center problems for the algebra and the corresponding space extending certain known results to a broader functional-analytic framework. The structure of the article is as follows: Section 2 provides the necessary background, including basic definitions, notations and preliminary results. Section 3 focuses on the existence of left and right identities in and explore their relationship with the space In particular, Theorem 3.4 characterizes the amenability of the group in terms of the existence of a right identity in under the assumption that the Young function satisfies the MA condition.
In Section 4, we prove Theorem 4.5, which provides a sufficient condition under which the restriction of an element in to lies in This result plays a central role in establishing the main theorem of the article regarding the topological center of the double dual of Section 5 is devoted to investigating the topological center problems for both the algebra and the corresponding space In Theorem 5.2, we establish a necessary and sufficient condition for the equality to hold, assuming that is amenable. This result can be seen as an Orlicz analogue of Miao’s theorem for algebras. Moreover, an analogous result (Theorem 5.4) is established for the topological center of the Banach algebra
2. Preliminaries and Some Notations
This section begins with a review of the fundamental terminology related to Orlicz spaces.
An even convex function is termed as a Young function if it satisfies the conditions and According to this definition, a Young function may attain the value at certain points and may therefore be discontinuous. However, in this context, we shall always assume that is real-valued and continuous.
For each Young function a corresponding convex function known as the complementary function to is defined by
The function thus defined is also a Young function. The pair (or equivalently ) is referred to as a complementary pair of Young functions, or a complementary Young pair. As an example, for the function is a Young function, and its complementary function is where and are dual Hölder exponents.
Let be a locally compact group equipped with a left Haar measure A Young function is said to satisfy the condition (globally), denoted , if there exists a constant and such that
for all [45, Definition 1, Pg. 22]. By [45, Remark, Pg. 46], is said to be regular if the condition holds locally for when the measure is finite and globally when the measure is infinite. A pair of complementary Young functions is said to satisfy the condition if both and individually satisfy the condition. It is straightforward to verify that the family of Young functions for and satisfies the condition. Another example is the function which satisfies the condition However, its complementary function does not satisfy the condition.
A Young function is said to satisfy the Milnes-Akimoni condition (abbreviated as the MA condition) if, for each there exists constants and such that
An example of a Young function satisfying the MA condition is for However, its complementary function for does not satisfy the MA condition. It is easy to verify that the family of Young functions defined by with satisfies the MA condition. Moreover, the corresponding family of complementary pairs of Young functions also satisfies the condition.
For a given Young function the associated Orlicz space is denoted by and is defined as
The space becomes a Banach space when equipped with the Luxemburg norm (also known as the Gauge norm), defined by
Let be the complementary function to Then, the Orlicz norm on is given by
In fact, these two norms are equivalent and for every the following holds:
If a Young function satisfies the condition, then the space of all continuous functions on with compact support is dense in Further, if the complementary function is continuous and satisfies if and only if then the dual space of is isometrically isomorphic to [45, Corollary 9, Pg. 111]. In particular, if the complementary Young pair satisfies the condition, then the Banach space is reflexive [45, Theorem 10, Pg. 112].
For more detailed information on Orlicz spaces, we refer the readers to [45].
Let be a Young function satisfying the condition. For any function , we define by for all The Orlicz Figà-Talamanca Herz algebra is defined as the set of all continuous functions that can be expressed as
where and such that
For any the norm is defined by
Equipped with this norm, along with pointwise addition and multiplication, the space becomes a commutative Banach algebra. Moreover, it is a regular, tauberian, semi-simple Banach algebra with its Gelfand spectrum homeomorphic to
Recall from [31, Lemma 4.1] that for any open subgroup of the Banach algebra is isometrically isomorphic to the subalgebra of consisting of functions that vanish outside We denote by the inclusion map from to and by the restriction map from to It follows from [31, Lemma 4.5] that the map is a contractive homomorphism of Banach algebras. Further, it is easy to verify that is a closed ideal of using the aforementioned identification. Moreover, the map is a bounded projection which is also a multiplier, i.e., it satisfies
for all and for all
Let denote the multiplier algebra of defined as the set
Each element induces a bounded linear operator given by for all Equipped with the operator norm and pointwise operations (addition and multiplication), the space is a commutative Banach algebra. Moreover, it is clear that
Let denote the set of all bounded complex Radon measures on For and we define the operator by left convolution:
If denote the Banach space of all bounded linear operators on equipped with the operator norm. It is straightforward to verify that for each Let be the closure of the set in with respect to the ultra-weak topology (the -topology). It is proved that the dual space of is isometrically isomorphic to [31].
Let denote the norm closure in of the set where is the measure associated with The dual of is denoted by which is a commutative Banach algebra containing [34]. For each the associated operator is denoted by The map thus defines the left regular representation of on
A linear functional is called a topologically invariant mean if it satisfies the following conditions:
for all and We denote the set of all topologically invariant means on by
For additional details on the Banach algebra and the related spaces, we refer the readers to the series of papers [1, 3, 7, 8, 9, 31, 32, 33, 34].
Let be a Banach algebra and let denote its double dual. It is well known that can be equipped with two natural extensions of the multiplication on both of which turn into a Banach algebra. These extensions, introduced by Arens [4], are commonly referred to as the first and second Arens product. In this article, we focus on the first Arens product, denoted by which is defined as follows:
Recall that a Banach algebra is said to be weakly sequentially complete (WSC) if every weakly Cauchy sequence converges with respect to the weak topology.
A locally compact group is said to be amenable if there exists a left-invariant mean on i.e., a positive linear functional on of norm one satisfying
where denotes the left translation of Classical examples of amenable groups include abelian groups, compact groups and solvable groups. In contrast, the free group on two generators is a well known example of a non-amenable group.
Throughout this article, denotes a locally compact group with a fixed Haar measure and assume that the complementary pair of Young functions satisfies the condition.
3. The Double Dual of and the Space
In this section, we present certain results related to the left and right identities in the double dual of when considered as a Banach algebra with the first Arens product. Theorem 3.4 establishes a necessary and sufficient condition for the existence of a right identity in Further, we explore their relationship with the space Moreover, the semi-simplicity of the Banach algebras and is studied in Theorem 3.6 and Theorem 3.9, respectively.
Remark 3.1.
We would like to mention that, to the best of our knowledge, the results presented in this section are, in fact, new for the Figà-Talamanca Herz algebra i.e., when the corresponding Young function is given by for
Let us begin with the following propositions related to the existence of left identity in the Banach algebra
Proposition 3.2.
The algebra has a left identity if is compact.
Proof.
Since is compact, the algebra contains the unit function 1 which also serves as the identity. Then, it is easy to verify by elementary calculations and using the definition of the Arens product that is a left identity for where is the canonical map from to ∎
Proposition 3.3.
If has a left identity, then
Proof.
If possible, assume that is a proper closed subspace of Then, by Hahn-Banach theorem, there exists a non-zero such that This implies that for all and hence, for all Since has a left identity, say we have,
which is a contradiction. Thus, ∎
Here is the main theorem of this section, which characterizes the amenability of the group in terms of the existence of a right identity in the algebra This result generalizes Lau’s result [35, Proposition 3.2 (a)] on for any locally compact group Furthermore, the analogous result for hypergroups is given by Esmailvandi et al. [15, Proposition 4.1].
Theorem 3.4.
Let satisfies the MA condition. Then the Banach algebra has a right identity if and only if is amenable.
Proof.
Assume that is amenable. Then, by [33, Theorem 3.1], the algebra possesses a bounded approximate identity. Hence, by [5, Proposition III.28.7], there exists a right identity in the Banach algebra
Conversely, let be a right identity in Now, by Goldstine’s theorem ([10, Theorem A.3.29]), there exists a net in such that
where is the canonical map from to
In the subsequent results, we investigate the semi-simplicity of the algebras and establishing certain necessary and sufficient conditions.
Proposition 3.5.
If the group is finite, then the algebra is semi-simple.
Proof.
The result follows from [7, Theorem 3.8] and the fact that the algebra is semi-simple. ∎
The following theorem demonstrates that the semi-simplicity of the algebra is sufficient for a group to be discrete, provided that is a second countable group for which the algebra is WSC. The precise statement of the result is as follows.
Theorem 3.6.
Let be a second countable group such that the algebra is weakly sequentially complete. If is semi-simple, then is discrete.
Proof.
Let
We claim that is an ideal in
Observe that, for any and
Since, the condition is equivalent to and the Arens product is associative, it follows that
This implies for all and
Further, by using the fact that the algebra is contained in the center of [9, Theorem 3.2], one can similarly prove that Thus, is an ideal in
Now, if possible, assume that and are two distinct elements of Then, by definition, which contradicts the fact that the Banach algebra is semi-simple. This implies that is singleton and hence, by [31, Corollary 7.6], is discrete. ∎
Proposition 3.7.
If is semi-simple, then
Proof.
Let us suppose to the contrary that is a proper closed subspace of Let be the annihilator of in i.e.,
Then, by Hahn-Banach theorem, Further, it is clear from the proof of Proposition 3.3 that is a left ideal in and In particular, we have, and hence, this contradicts the semi-simplicity of the algebra Thus, ∎
Remark 3.8.
Note that the Banach algebras and are semi-simple. This follows directly from the relation [34, Corollary 3.2] and the fact that the set is contained in the Gelfand spectrum of all of these algebras.
The following theorem addresses the semi-simplicity of the algebra and is analogous to Theorem 3.6.
Theorem 3.9.
The Banach algebra is semi-simple if is discrete. The converse holds when is a second countable group for which is weakly sequentially complete.
Proof.
Since is discrete, by [32, Corollary 4.13 (ii)], and thus, [34]. The desired result follows directly from Remark 3.8.
Conversely, let be a second countable group such that the algebra is weakly sequentially complete and assume that is semi-simple. Consider
Now, by repeating the arguments as in Theorem 3.6, it follows that is an ideal in and Since, by [31, Corollary 6.2], let be the restriction of to for some non-zero Observe that is also non-zero and belongs to Otherwise, if then
As
Now, by choosing a suitable such that we have,
This contradicts the fact that is non-zero and hence, Thus, by repeating the steps as in Theorem 3.6, it follows that is a singleton set and hence, by [31, Corollary 7.6], is discrete. ∎
4. Restriction Theorem
The aim of this section is to prove the restriction theorem (Theorem 4.5), which plays a pivotal role in establishing the main result of this article (Theorem 5.2), concerning the topological center of the double dual of The ideas behind the proofs are motivated by [44].
We begin with the following proposition.
Proposition 4.1.
Let be an open subgroup of an amenable group Then every right unit of in admits an extension to a right unit of
Proof.
Recall from [29, Pg. 101] that for an algebra of functions the support of a linear functional as a subset of is characterized as follows: if and only if there exists a neighborhood of such that for all with
Remark 4.2.
It is easy to verify that if for and then the function
Lemma 4.3.
Let be an open subgroup of and be the restriction map. Then for all with
Proof.
Let be such that and for all with Then, for any with we have,
Now, by Remark 4.2, it follows that
Thus, by the density of in it follows that for all with ∎
The next lemma is a crucial component in proving Theorem 4.5 and establishes the existence of a sequence of compact subsets of with certain significant properties.
Lemma 4.4.
Let and assume that for every open compact subgroup of is in Then, for each there exists a compact subset of such that for all with compact support, and
Proof.
Let us suppose to the contrary that there exists and a function with compact support such that and where Now, by reiterating the construction outlined in [44, Lemma 3.1], we obtain a sequence of compactly supported functions in and a sequence of relatively compact symmetric neighborhoods of in such that, for each
and
and
Let Then, is an open compact subgroup of and by hypothesis, is in This implies that for some where is the canonical map from to Now, we claim that there exists a compact subset of such that for any with and Since is dense in there exists such that
Let be a relatively compact neighborhood of in Then, the required compact set is and observe that, for any with and we have,
In the above computation, since the term equals zero [29, Pg. 101]. Now,
i.e., is an open cover of Since, is compact and the sequence is increasing, there exists such that Then,
and thus, by the above claim, it follows that However, by Lemma 4.3, we have,
which is a contradiction. Hence, the result follows. ∎
Here is the main theorem of this section, which provides a sufficient condition for any under which the restriction of to belongs to The analogous result for the Fourier algebra where is a hypergroup and for Figà-Talamanca Herz algebra where is a locally compact group, are given by Esmailvandi et al. [15, Proposition 5.5] and Miao [44, Lemma 3.1], respectively. The following theorem is the Orlicz analogue of Miao’s result.
Theorem 4.5.
Let If for every open compact subgroup of is in then the restriction of to belongs to
Proof.
Let be an open compact subgroup of containing all compact subsets of from Lemma 4.4. Then, for any with compact support, it is easy to verify that the function also belongs to and has compact support contained in Since, for each by Lemma 4.4, it follows that Thus,
for all with compact support and hence, by [32, Corollary 4.7], the restriction of to is The desired result is a direct consequence of the hypothesis that is in ∎
5. Topological center
In this section, we study the topological center of the double dual of when considered as a Banach algebra with the first Arens product. The primary aim is to prove Theorem 5.2, which establishes a necessary and sufficient condition for the equality to hold, assuming that is amenable. The ideas and results are inspired by Miao’s work on algebras [44].
Let us begin with a straightforward yet significant lemma.
Lemma 5.1.
Let be an open subgroup of and be the inclusion map. Then the following holds
Proof.
Since the restriction map is a bounded projection which is also a multiplier, the result is a direct consequence of [44, Theorem 4.2]. ∎
Here is the main result of this article concerning the topological center of the double dual of This theorem serves as an Orlicz analogue of Miao’s result on algebras [44, Theorem 4.4].
Theorem 5.2.
Let be an amenable group and suppose that satisfies the MA condition. Then if and only if for every open compact subgroup of
Proof.
Assume that Let be an open compact subgroup of and Then, by Lemma 5.1, we have,
So, Now, by elementary calculations, it is easy to verify that the function vanishes outside and thus, by the usual identification of in it follows that belongs to Hence,
Conversely, assume that for any open compact subgroup of and let Then, by Lemma 5.1, we have,
So, Now, by Theorem 4.5, the restriction of (say ) to is in i.e., for some where is the canonical map from to Since is amenable, by [33, Theorem 3.1], the algebra has a bounded approximate identity, say and further, by Theorem 3.4, the algebra has a right identity, say Moreover, it is clear from the proof of [5, Proposition III.28.7] that is the cluster point of Thus, for any since the function is in we have,
Furthermore, since converges in topology to and it follows that,
This implies that i.e., is in Hence, ∎
In the subsequent results, we study the topological center of the Banach algebra From [8, Corollary 3.4], we recall that the space is continuously embedded in under the assumption that the group is amenable. Further, it is clear from the identification in [8, Theorem 3.2 and Corollary 3.4] that for any the element in is given by
where and
We now present the following important lemma.
Lemma 5.3.
For any amenable group the space is contained in the topological center of i.e.,
Proof.
To prove this lemma, it is enough to show that the space is contained in the center of as in [36, Proposition 4.5]. Now, let and Then, for any and we have,
On the other hand,
Thus,
for all and Hence, the result follows. ∎
The final theorem of the article deals with the topological center of the algebra and is analogous to Theorem 5.2. The proof of the theorem follows verbatim to [44, Corollary 4.6 (ii)] and hence, we omit it. The main ingredients of the proof are Lemma 5.3 and the following two facts:
Every closed subgroup of an amenable group is amenable.
For an amenable group the space which is the predual of is contained in This follows from [32, Corollary 4.13 (i)], since is amenable.
Here is the exact statement of the theorem.
Theorem 5.4.
Let be an amenable group. Then if and only if for every open compact subgroup of
Acknowledgement
The first author would like to thank the Indian Institute of Technology Delhi for providing the Institute Assistantship.
Data Availability
Data sharing does not apply to this article as no datasets were generated or analysed during the current study.
Competing Interests
The authors declare that they have no competing interests.